OpenStudy (howard-wolowitz):

Use the information given to identify the .... term of the geometric sequence:

OpenStudy (howard-wolowitz):

\[a _{9}\]

OpenStudy (howard-wolowitz):

\[a _{9}=5 ...... r=-2\]

OpenStudy (howard-wolowitz):

\[a _{9}= 2,560...-320...320...-2,560\]

OpenStudy (howard-wolowitz):

@pitamar

OpenStudy (anonymous):

Geometric sequences are of the form: $$a_n = c \cdot r^n$$ Where \(c\) and \(r\) are constants. We are not told \(c\) but we're told that \(r=-2\). So we have: $$a_n = c \cdot (-2)^n$$ Now we are told that \(a_9 = 5\). But we know that: $$ a_5 = c \cdot (-2)^9$$ Which means: $$ 5 = c \cdot (-2)^9 $$So first step is to find \(c\). Can you do that?

OpenStudy (howard-wolowitz):

yeah to get?

OpenStudy (howard-wolowitz):

-2,560

OpenStudy (howard-wolowitz):

because of those steps

OpenStudy (howard-wolowitz):

actually i made a error sorry

OpenStudy (anonymous):

I am actually sorta confused. I see that in your question you are already told that \(a_9 = 5\)... so what are you supposed to find?

OpenStudy (howard-wolowitz):

what a9 equals

OpenStudy (anonymous):

but you are told \(a_9 = 5\) at the beginning. Is there a mistake there?

OpenStudy (howard-wolowitz):

no thats supposed to be the equation

OpenStudy (anonymous):

Ok, I'm confused. If you're told \(a_9\) then why are you supposed to find it? sorry about that.

OpenStudy (howard-wolowitz):

you use it for something but no im confused

OpenStudy (howard-wolowitz):

@SolomonZelman

OpenStudy (howard-wolowitz):

@welshfella

OpenStudy (howard-wolowitz):

its saying use that to find it

OpenStudy (howard-wolowitz):

dont leaveeeee

OpenStudy (howard-wolowitz):

so what is c?

OpenStudy (anonymous):

Well, though the question now seems pointless.... C is the constant that will make $$ 5 = c \cdot (-2)^9 $$ Work. We just have to solve the equation. So we divide both sides by (-2)^9: $$ \frac{5}{(-2)^9} = c $$And since \((-2)^9 = -512\): $$ c = -\frac{5}{512} \\ a_n = -\frac{5}{512} \cdot (-2)^n $$ But it makes no sense in the question.. we got this because we know what \(a_9\) was... can you get a picture of this question?

OpenStudy (howard-wolowitz):

how would i do that?

OpenStudy (howard-wolowitz):

like how do i do that?

OpenStudy (anonymous):

Is it from a book or website?

OpenStudy (howard-wolowitz):

a website

OpenStudy (anonymous):

http://www.take-a-screenshot.org/

OpenStudy (anonymous):

When you save it in Paint, make sure you save it in format .jpg (also called .jpeg) or .png Do not save as .bmp (bitmap), it's really heavy...

OpenStudy (howard-wolowitz):

i took it i think so now what

OpenStudy (howard-wolowitz):

i dont have paint

OpenStudy (anonymous):

How comes? you use mac?

OpenStudy (anonymous):

When you have the picture file you can upload it here. There is an 'Attach File' button down here. What is your operating system then?

OpenStudy (howard-wolowitz):

when i take the picture where does it go so i can attach it???

OpenStudy (anonymous):

It goes to clipboard. you can paste it in programs. Like Paint or Office Word. Do you have any of these?

OpenStudy (howard-wolowitz):

yes but wheres the clipboard

OpenStudy (howard-wolowitz):

i just cant find the clipboard

OpenStudy (anonymous):

Oh lol. Clipboard is just a name for things you've copied. You know Copy Paste? Ctrl-C Ctrl-V? So it's like you 'copied' the picture. Now just paste it.

OpenStudy (howard-wolowitz):

ok

OpenStudy (howard-wolowitz):

like that?

OpenStudy (anonymous):

Yes, very good. Man... I think this is a mistake in the question lol =\ It makes absolutely no sense...

OpenStudy (howard-wolowitz):

well what would you guees the answer is cause i have no idea

OpenStudy (anonymous):

Sorry the site has stopped working for me. I think their mistake is that they wrote \(a_9 = 5 \qquad r=-2\) but meant \(c = 5 \qquad r=-2\), because in that case our equation becomes: $$ a_n = c \cdot r^n = 5 \cdot (-2)^n $$And: $$ a_9 = 5 \cdot (-2)^9 = 5 \cdot -512 = -2560 $$Which is one of the answers.